OpenStudy (anonymous):

Which value of x satisfies both -9x + 4y = 8 and -3x − y = 4?

OpenStudy (gebooors):

-9x + 4y = 8 - 3x - y = 4 Multiply lower equation (for instance) 4 and You eliminate y when you add equations together. Then solve x ja then y.